# Out-of-equlibrium inference of feeding rates through population data from generic consumer-resource stochastic dynamics

### v2

@@ Abstract
Statistical models are often structurally unidentifiable, because different set of parameters can lead to equal model outcomes. To be useful for prediction and parameter inference from data, stochastic population models need to be identifiable, this meaning that model parameters can be uniquely inferred from a large number of model observations. In particular, precise estimation of feeding rates in consumer-resource dynamics is crucial, because consumer-resource processes are central in determining biomass transport across ecosystems. Model parameters are usually estimated at stationarity, because in that case model analyses are often easier. In this contribution we analyze the problem of parameter redundancy in a multi-resource consumer-resource model, showing that model indentifiability depends on whether the dynamics have reached stationarity or not. To be precise, we: (i) Calculate the steady-state and out-of-equilibrium probability distributions of predator’s abundances analytically using generating functions, which allow us to unveil parameter redundancy and carry out proper maximum likelihood estimation. (ii) Conduct in silico experiments by tracking the abundance of consumers that are either searching for or handling prey, data then used for maximum likelihood parameter estimation. (iii) Show that, when model observations are recorded out of equilibrium, feeding parameters are truly identifiable, whereas if sampling is done at stationarity, only ratios of rates can be inferred from data (i.e., parameters are redundant). We discuss the implications of our results when inferring parameters of general dynamical models.
@@ 1 Introduction
Stochastic models are to be designed so that model parameters can be uniquely determined from a large number of model realizations. In that case, the model is known as structurally identifiable, because model parameters can be inferred without redundancy from data, regardless the quantity or quality of the data (Bellman and Åström, 1970). In particular, dynamic consumer-resource models should allow per-capita feeding rates (associated to elemental feeding processes) to be estimated unambiguously from experimental population abundance data. The problem of structural identifiability has been thoroughly analyzed for deterministic dynamics in the context of general biological modeling approaches (see Villaverde et al. (2016) and references therein). In this contribution, we focus on the problem of structural identifiability in stochastic dynamical models. Our aim is to unveil scenarios for feeding rate redundancy using a multi-resource, Holling type II feeding dynamics (Holling, 1959a, b). For that purpose, we consider a single predator population, in which individuals are able to search for and handle $S$ different resource species, leading to $S+1$ different behavioral types, namely: $S$ types associated to predators handling each resource species, and free predators that are not handling any resource. We assume mass action and chemostatic conditions to derive the joint probability distribution of the different behavioral types, at two regimes: in the transient dynamics (out of equilibrium) and in the long run (once the steady-state is reached). We show analytically that, at equilibrium, the model is not identifiable because only the ratio between relaxation and attack rates can be extracted from fitting model to data. In this sense, the model at equilibrium is set identifiable, this meaning that parameters can be precisely determined only on a subset of the parameter space. We also show that true feeding rates cannot be properly estimated via a standard fitting procedure such as maximum likelihood (ML) parameter estimation using synthetic (model-simulated) data at stationarity. This in silico experiment, which mimics an empirical setup to estimate feeding parameters that assumes the system has reached stationarity, only allows for precise estimation of quotients of rates at equilibrium.
@@ 2 Stochastic model identifiability via generating functions
$\mathcal{P}=\{P_{\bm{\theta}}|\bm{\theta}\in\Theta\}.$ (1)
@@
For an identifiable model, the map $\bm{\theta}\mapsto P_{\bm{\theta}}$ can be inverted to learn the true parameter set if the model can be observed an infinite number of times. This is true because of the strong law of large numbers, by which probabilities can be estimated by averaging frequencies over a large number of observations (Loève and Loève, 1977). Thus, with an infinite number of observations, finding the true probability distribution $P_{\bm{\theta}_{0}}$ is feasible. Because the map $\bm{\theta}\mapsto P_{\bm{\theta}}$ is invertible, the true parameter set $\bm{\theta}_{0}$ which generated the distribution $P_{\bm{\theta}_{0}}$ can be unambiguosly extracted from the distribution.
@@
$G(z,t)=\frac{\left(\frac{\beta}{\delta}-1\right)^{\frac{\mu}{\beta}}\left[e^{(\beta-\delta)t}(1-z)+\frac{\beta}{\delta}z-1\right]^{n_{0}}}{\left[\frac{\beta}{\delta}e^{(\beta-\delta)t}(1-z)+\frac{\beta}{\delta}z-1\right]^{n_{0}+\frac{\mu}{\beta}}}$ (7)
@@ 3 Multi-resource Holling Type II feeding dynamics
$T(\bm{n}+\bm{e}_{i}|\bm{n})=\alpha_{i}\frac{n_{A}n_{R_{i}}}{N},\qquad T(\bm{n}-\bm{e}_{i}|\bm{n})=\nu_{i}n_{i},$ (12)
@@
Our aim is to estimate properly feeding rates $\{\alpha_{1},\dots,\alpha_{S},\nu_{1},\dots,\nu_{S}\}$. We first compute the steady-state distribution for arbitrary parameter values, and show that the distribution of feeding types is multinomial. To do that, we use a detailed-balance condition satisfied by transition probabilities. Second, using the generating function formalism, we find an analytical solution for out-of-equilibrium probabilities $P_{\bm{n}}(t)$ for the particular case of equal relaxation rates, $\nu_{1}=\dots=\nu_{S}$, and for an initial state formed only by searching predators. In the second case, we also show that the distribution is multinomial. From the form of the distributions (and the generating function), it is apparent that the model is perfectly identifiable out of equilibrium.
3.1 Steady-state distribution
$P_{n_{A}}=\binom{n_{A}^{0}}{n_{A}}p^{n_{A}}(1-p)^{n_{A}^{0}-n_{A}},$ (15)
$n_{A}$ being the number of searching consumers (Palamara et al., 2022). One can take advantage of knowing that distribution to rigorously derive the functional response associated to the single reaource case ($S=1$), which turns out to be the Holling type II functional response (Palamara et al., 2022). We here extend the binomial distribution result to a multi-resource case by solving the master equation at stationarity, which yields an analytical expression for the joint steady-state distribution $P_{\bm{n}}$. We can do this because transition probability rates (13) are linear in population numbers (Kendall, 1949; Nisbet and Gurney, 2004). As a result, the single-resource binomial distribution is extended to a multi-resource case, leading to a multinomial distribution for the joint probability distribution of observing a vector $(n_{A},\bm{n})$ of the populations of behavioral types, as we show next.
$T(\bm{n}-\bm{e_{i}}|\bm{n})P_{\bm{n}}=T(\bm{n}|\bm{n}-\bm{e_{i}})P_{\bm{n}-\bm{e_{i}}},$ (18)
and, subsequently,
$P_{\bm{n}+\bm{e}_{i}}=\frac{\nu_{j}}{\theta_{j}}\frac{\theta_{i}}{\nu_{i}}\frac{n_{j}+1}{n_{i}+1}P_{\bm{n}+\bm{e}_{j}}.$ (20)
Making the change $n_{i}\to n_{i}-1$ in the last equation one gets
Now, apply Eq. (21) to resources $i=1$ and $j=2$: this gives
$P_{\bm{n}}=\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\frac{n_{2}+1}{n_{1}}P_{(n_{1}-1,n_{2}+1,\dots,n_{S})}.$ (22)
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\right)^{2}\frac{(n_{2}+1)(n_{2}+2)}{n_{1}(n_{1}-1)}P_{(n_{1}-2,n_{2}+2,\dots,n_{S})}.$ (23)
By repeating the iteration $n_{1}-2$ additional times one gets
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\right)^{n_{1}}\binom{n_{1}+n_{2}}{n_{1}}P_{(0,n_{1}+n_{2},\dots,n_{S})},$ (24)
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\right)^{n_{1}}\left(\frac{\theta_{2}}{\nu_{2}}\frac{\nu_{3}}{\theta_{3}}\right)^{n_{1}+n_{2}}\binom{n_{1}+n_{2}}{n_{1}}\binom{n_{1}+n_{2}+n_{3}}{n_{1}+n_{2}}P_{(0,0,n_{1}+n_{2}+n_{3},\dots,n_{S})},$ (25)
which simplifies to
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\right)^{n_{1}}\left(\frac{\theta_{2}}{\nu_{2}}\right)^{n_{2}}\left(\frac{\nu_{3}}{\theta_{3}}\right)^{n_{1}+n_{2}}\binom{n_{1}+n_{2}+n_{3}}{n_{1},n_{2},n_{3}}P_{(0,0,n_{1}+n_{2}+n_{3},\dots,n_{S})},$ (26)
$P_{\bm{n}-\bm{e}_{S}}=\frac{\nu_{S}n_{S}}{\theta_{1}(n_{A}^{0}+1-\sum_{k=1}^{S}n_{k})}P_{\bm{n}}.$ (28)
$P_{(q-1)\bm{e}_{S}}$ $=\frac{\nu_{S}}{\theta_{S}}\frac{q}{n_{A}^{0}+1-q}P_{q\bm{e}_{S}},$ (29)
[…]
$P_{q\bm{e}_{S}}=\binom{n_{A}^{0}}{q}\left(\frac{\theta_{S}}{\nu_{S}}\right)^{q}P_{\bm{0}}.$ (30)
where we have used Eq. (11) and
$\binom{n_{A}^{0}}{\sum_{k=1}^{S}n_{k}}\binom{\sum_{k=1}^{S}n_{k}}{n_{1},\dots,n_{S}}=\frac{n_{A}^{0}!}{n_{1}!\cdots n_{S}!n_{A}!}.$ (32)
In the form of Eq. (31), the joint stationary distribution remarkably resembles a multinomial distribution $\mathcal{M}(n_{A}^{0},\{p_{k}\})$ for random variables $(X_{1},\dots,X_{S})$, whose density function is defined in general as
$P_{\bm{n}}=\begin{cases}\frac{n_{A}^{0}!}{\prod_{k=1}^{S}n_{k}!}\prod_{k=1}^{S}p_{k}^{n_{k}},&\text{when }\sum_{k=1}^{S}n_{k}=n_{A}^{0},\\ 0,&\text{otherwise},\end{cases}$ (33)
the distribution being determined by the number of trials $n_{A}^{0}$ and a set of probability parameters $\{p_{k}\}$ which satisfy $\sum_{k=1}^{S}p_{k}=1$. To write Eq. (31) as a multinomial distribution, first we have to calculate the normalization constant $P_{\bm{0}}$. It follows immediately after evaluating the probability of observing a total number $q$ of handling predators, $q=\sum_{k=1}^{S}n_{k}$, which can be done as follows.
$P_{q}=\sum_{\sum_{i}n_{i}=q}\frac{n_{A}^{0}!}{n_{1}!\cdots n_{S}!n_{A}!}\left[\prod_{k=1}^{S}\left(\frac{\theta_{k}}{\nu_{k}}\right)^{n_{k}}\right]P_{\bm{0}},$ (34)
which can be written alternatively as
$P_{q}=\binom{n_{A}^{0}}{q}\sum_{\sum_{i}n_{i}=q}\frac{q!}{n_{1}!\cdots n_{S}!}\left[\prod_{k=1}^{S}\left(\frac{\theta_{k}}{\nu_{k}}\right)^{n_{k}}\right]P_{\bm{0}}.$ (35)
$\sum_{\sum_{i}n_{i}=q}\frac{q!}{n_{1}!\cdots n_{S}!}\prod_{k=1}^{S}\left(\frac{\theta_{k}}{\nu_{k}}\right)^{n_{k}}=\left(\sum_{k=1}^{S}\frac{\theta_{k}}{\nu_{k}}\right)^{q},$ (36)
which implies that
Therefore, if we impose normalization of the distribution $P_{q}$ we get
where in the last equality we have used the binomial theorem.
The joint distribution of the vector of population abundances of different behavioral types, $(n_{A},n_{1},\dots,n_{S})$, is given by Eqs. (31) and (38). It is a multinomial distribution $\mathcal{M}(n_{A}^{0},\{p_{A},p_{1},\dots,p_{S}\})$, as defined in Eq. (33), with $n_{A}^{0}$ trials and probabilities given by
@@ 3.1 Steady-state distribution
for predators handling on resources $i=1,\dots,S$. In addition,
$p_{A}=\frac{1}{1+\sum_{k=1}^{S}\frac{\theta_{k}}{\nu_{k}}}=1-\sum_{k=1}^{S}p_{k}$ (40)
is the probability associated to the number of free predators $n_{A}$. Notice that Eq. (40) extends naturally, for more than one resource, the probability parameter $p=\left(1+\frac{\theta}{\nu}\right)^{-1}$ of the binomial distribution (15).
and a mean number of free predators equal to
In addition to the joint, steady-state probability distribution of different behavioral types, by summing over the corresponding variables we can obtain the marginal distributions of each single type, which are binomials $B(n_{A}^{0},p_{i})$ (Loève and Loève, 1977). In particular, from our derivation above we find that the distribution of the total number of handling consumers is a binomial $B(n_{A}^{0},1-p_{A})$: using Eq. (38) and (40), the probability that $\sum_{k=1}^{S}n_{k}=q$, which is given by Eq. (37), can be rewritten as
$P_{q}=\binom{n_{A}^{0}}{q}(1-p_{A})^{q}p_{A}^{n_{A}^{0}-q}.$ (43)
$P_{n_{A}}=\binom{n_{A}^{0}}{n_{A}}p_{A}^{n_{A}}(1-p_{A})^{n_{A}^{0}-n_{A}},$ (44)
as expected (Palamara et al., 2022).
@@ 3.2 Out-of-equilibrium probability distribution
$\frac{\partial G}{\partial t}=\sum_{i=1}^{S}(1-z_{i})\biggl{(}\nu_{i}\frac{\partial G}{\partial z_{i}}+\theta_{i}\sum_{j=1}^{S}z_{j}\frac{\partial G}{\partial z_{j}}-\theta_{i}n_{A}^{0}G\biggr{)}.$ (45)
@@
We leave the details of the derivation of Eq. (46) from (14) to A.
@@
$K=\frac{\nu+R_{0}}{Q-R_{0}}=\frac{\nu+\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}}{\sum_{j=1}^{S}\theta_{j}(1-\hat{z}_{j})},$ (57)
where we have defined $R_{0}:=\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}$. Observe that the initial condition vector $\bm{\hat{z}}$ depends on the vector of $z$-variables, $\bm{z}$, via the explicit solution of the differential equations for $z_{i}$ in Eq. (52). Because the solution of the system of ODEs satisfied by the characteristic curves will depend on the vector $\bm{\hat{z}}$ of initial conditions, we have to substitute $\hat{z}_{j}$ as function of $\bm{z}$ at the end of the calculation in order to get the desired expression for the generating function $G(\bm{z},t)$.
@@
$R=\frac{KQ-\nu e^{(\nu+Q)t}+e^{\nu t}[(1+K)R_{0}+\nu-KQ]}{K+e^{(\nu+Q)t}},$ (62)
where we have defined $R=R(\bm{z}):=\sum_{j=1}^{S}\theta_{j}z_{j}$. Because $\hat{z}_{j}$ are at the end functions of $\bm{z}$, $K$ becomes itself a function of $\bm{z}$, which can be determined as follows. From Eq. (57) we can solve for $R_{0}=\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}$ in terms of $K$ as $R_{0}=(KQ-\nu)/(1+K)$ which, upon substitution into the last expression, yields
$R=\frac{KQ-\nu e^{(\nu+Q)t}}{K+e^{(\nu+Q)t}}.$ (63)
In this equality, any dependence on the initial condition $\bm{\hat{z}}$ has been removed. We can solve for $K$ in terms of $R$, giving
$K(\bm{z})=\frac{R(\bm{z})+\nu}{Q-R(\bm{z})}e^{(\nu+Q)t}.$ (64)
Finally, an explicit expression of initial conditions $\hat{z}_{j}$ as function of $\bm{z}$ is determined by substituting $K(\bm{z})$, given by Eq. (64), into (61), which results in an equation from which we can eliminate $\hat{z}_{j}$ in terms of $\bm{z}$. This manipulation gives
$\hat{z}_{j}=1-\left(1+\frac{\nu}{Q}\right)\frac{Q-R(\bm{z})+[R(\bm{z})-Qz_{j}]e^{Qt}}{Q-R(\bm{z})+[R(\bm{z})+\nu]e^{(\nu+Q)t}}.$ (65)
@@
$G(\bm{z},t)=\left[\frac{R({\bm{z}})+\nu+(Q-R({\bm{z}}))e^{-(\nu+Q)t}}{\nu+Q}\right]^{n_{A}^{0}}\prod_{j=1}^{S}\left[1-\left(1+\frac{\nu}{Q}\right)\frac{Q-R(\bm{z})+(R(\bm{z})-Qz_{j})e^{Qt}}{Q-R(\bm{z})+(R(\bm{z})+\nu)e^{(\nu+Q)t}}\right]^{\hat{n}_{j}}.$ (66)
Recall the definitions $R(\bm{z}):=\sum_{j=1}^{S}\theta_{j}z_{j}$ and $Q:=\sum_{j=1}^{S}\theta_{j}$, and that the condition $n_{A}^{0}-\sum_{j=0}^{S}\hat{n}_{j}\geq 0$ must hold for the last expression to make sense. This function satisfies the initial condition (49) as well as the normalization condition $G(\bm{1},t)=1$. The generating function has to be expanded in power series of $z_{i}$ in order to obtain the explicit expression of the probability distribution $P_{\bm{n}}(t)$. Such task is hard in general. As shown by the derivation itself, it is very difficult to find the out-of-equilibrium distribution from the master equation, in contrast with the solution at stationarity: for the former, we cannot find it unless assuming equal handling rates, and we cannot carry out the series expansion to obtain the probability distribution $P_{\bm{n}}(t)$ for arbitrary initial conditions.
@@
$p_{0}(t)$ $=\frac{1+e^{-\left(\nu+\sum_{k=1}^{S}\theta_{k}\right)t}\sum_{i=1}^{S}\frac{\theta_{i}}{\nu}}{1+\sum_{k=1}^{S}\frac{\theta_{k}}{\nu}},$ (68)
$p_{j}(t)$ $=\frac{\theta_{j}}{\nu}\frac{1-e^{-\left(\nu+\sum_{k=1}^{S}\theta_{k}\right)t}}{1+\sum_{k=1}^{S}\frac{\theta_{k}}{\nu}},$
for $j=1,\dots,S$. It is easy to check that $0\leq p_{j}(t)\leq 1$ for $j=0,1,\dots,S$ and arbitrary time $t$. In addition, it holds that $\sum_{j=0}^{S}p_{j}(t)=1$ for any time. Therefore, the parameter combinations $p_{j}(t)$ can be interpreted as probabilities. As we immediately show below, they are precisely the probabilities associated to the $j$-th population in a multinomial distribution.
@@
$G(\bm{z},t)=\biggl{(}p_{0}(t)+\sum_{j=1}^{S}p_{j}(t)z_{j}\biggr{)}^{n_{A}^{0}}=\sum_{n_{A}=0}^{n_{A}^{0}}\binom{n_{A}^{0}}{n_{A}}p_{0}^{n_{A}}\biggl{(}\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}-n_{A}}.$ (69)
@@
$\biggl{(}\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}-n_{A}}=\!\!\!\!\sum_{\begin{subarray}{c}n_{1},\dots,n_{S}\geq 0\\ \sum_{i}n_{i}=n_{A}^{0}-n_{A}\end{subarray}}\!\!\!\!\frac{(n_{A}^{0}-n_{A})!}{n_{1}!\cdots n_{S}!}\prod_{k=1}^{S}(p_{k}z_{k})^{n_{k}},$ (70)
@@
$P_{\bm{n}}(t)=\sum_{\begin{subarray}{c}n_{A}=0\\ n_{A}=n_{A}^{0}-\sum_{i}n_{i}\end{subarray}}^{n_{A}^{0}}\frac{n_{A}^{0}!}{n_{A}!n_{1}!\cdots n_{S}!}p_{0}^{n_{A}}p_{1}^{n_{1}}\cdots p_{S}^{n_{S}}$ (71)

### v3

@@ Abstract
Statistical models are often structurally unidentifiable, because different sets of parameters can lead to equal model outcomes. To be useful for prediction and parameter inference from data, stochastic population models need to be identifiable, this meaning that model parameters can be uniquely inferred from a large number of model observations. In particular, precise estimation of feeding rates in consumer-resource dynamics is crucial, because consumer-resource processes are central in determining biomass transport across ecosystems. Model parameters are usually estimated at stationarity, because in that case model analyses are often easier. In this contribution we analyze the problem of parameter redundancy in a multi-resource consumer-resource model, showing that model indentifiability depends on whether the dynamics have reached stationarity or not. To be precise, we: (i) Calculate the steady-state and out-of-equilibrium probability distributions of predator’s abundances analytically using generating functions, which allow us to unveil parameter redundancy and carry out proper maximum likelihood estimation. (ii) Conduct in silico experiments by tracking the abundance of consumers that are either searching for or handling prey, data then used for maximum likelihood parameter estimation. (iii) Show that, when model observations are recorded out of equilibrium, feeding parameters are truly identifiable, whereas if sampling is done solely at stationarity, only ratios of rates can be inferred from data (i.e., parameters are redundant). We discuss the implications of our results when inferring parameters of general dynamical models.
@@ 1 Introduction
Stochastic models are to be designed so that model parameters can be uniquely determined from a large number of model realizations. In that case, the model is known as structurally identifiable, because model parameters can be inferred without redundancy from data, regardless the quantity or quality of the data (Bellman and Åström, 1970). In particular, dynamic consumer-resource models should allow per-capita feeding rates (associated to elemental feeding processes) to be estimated unambiguously from experimental population abundance data. The problem of structural identifiability is general and of multidisciplinar character. For example, it has been thoroughly analyzed for deterministic dynamics in the context of biological modeling approaches (see Villaverde et al. (2016) and references therein), but also in social sciences (Helbing et al., 2015; Jusup et al., 2022). In this contribution, we focus on the problem of structural identifiability in stochastic dynamical models. Our aim is to unveil scenarios for feeding rate redundancy using a multi-resource, Holling type II feeding dynamics (Holling, 1959a, b). For that purpose, we consider a single predator population, in which individuals are able to search for and handle $S$ different resource species, leading to $S+1$ different behavioral types, namely: $S$ types associated to predators handling each resource species, and free predators that are not handling any resource. We assume mass action and chemostatic conditions to derive the joint probability distribution of the different behavioral types, at two regimes: in the transient dynamics (out of equilibrium) and in the long run (once the steady-state is reached). We show analytically that, at equilibrium, the model is not identifiable because only the ratio between relaxation and attack rates can be extracted from fitting model to data. In this sense, the model at equilibrium is set identifiable, this meaning that parameters can be precisely determined only on a subset of the parameter space. We also show that true feeding rates cannot be properly estimated via a standard fitting procedure such as maximum likelihood (ML) parameter estimation using synthetic (model-simulated) data at stationarity. This in silico experiment, which mimics an empirical setup to estimate feeding parameters that assumes the system has reached stationarity, only allows for precise estimation of quotients of rates at equilibrium.
@@ 2 Stochastic model identifiability via generating functions
$\mathcal{P}=\{P_{\bm{\theta}}|\bm{\theta}\in\Theta\}.$
@@
For an identifiable model, the map $\bm{\theta}\mapsto P_{\bm{\theta}}$ can be inverted to learn the true parameter set if the model can be observed an infinite number of times. This is true because of the strong law of large numbers, by which probabilities can be estimated by averaging frequencies over a large number of observations (Loève, 1977). Thus, with an infinite number of observations, finding the true probability distribution $P_{\bm{\theta}_{0}}$ is feasible. Because the map $\bm{\theta}\mapsto P_{\bm{\theta}}$ is invertible, the true parameter set $\bm{\theta}_{0}$ which generated the distribution $P_{\bm{\theta}_{0}}$ can be unambiguosly extracted from the distribution.
@@
$G(z,t)=\frac{\left(\frac{\beta}{\delta}-1\right)^{\frac{\mu}{\beta}}\left[e^{(\beta-\delta)t}(1-z)+\frac{\beta}{\delta}z-1\right]^{n_{0}}}{\left[\frac{\beta}{\delta}e^{(\beta-\delta)t}(1-z)+\frac{\beta}{\delta}z-1\right]^{n_{0}+\frac{\mu}{\beta}}}$
@@ 3 Multi-resource Holling Type II feeding dynamics
$T(\bm{n}+\bm{e}_{i}|\bm{n})=\alpha_{i}\frac{n_{A}n_{R_{i}}}{N},\qquad T(\bm{n}-\bm{e}_{i}|\bm{n})=\nu_{i}n_{i},$
@@
Our aim is to accurately estimate feeding rates $\{\alpha_{1},\dots,\alpha_{S},\nu_{1},\dots,\nu_{S}\}$. We first compute the steady-state distribution for arbitrary parameter values, and show that the distribution of feeding types is a multinomial $\mathcal{M}(n_{A}^{0},\{p_{A},p_{1},\dots,p_{S}\})$ with index $n_{A}^{0}$ and probability parameters $\{p_{A},p_{1},\dots,p_{S}\}$. Labeling $p_{0}:=p_{A}$, the multinomial density function is defined as
$P_{\bm{n}}=\begin{cases}\frac{n_{A}^{0}!}{\prod_{k=0}^{S}n_{k}!}\prod_{k=0}^{S}p_{k}^{n_{k}},&\text{when }n_{A}+\sum_{k=1}^{S}n_{k}=n_{A}^{0},\\ 0,&\text{otherwise}.\end{cases}$ (11)
Here, vector $\bm{n}=(n_{A},n_{1},\dots,n_{S})$ gathers the number of searching predators, $n_{A}$, together with the numbers $n_{i}$ of predators handling on each resource ($i=1,\dots,S$). Probability parameters are
@@
for $i=1,\dots,S$, and
$p_{A}=p_{0}=\frac{1}{1+\sum_{k=1}^{S}\frac{\theta_{k}}{\nu_{k}}}=1-\sum_{k=1}^{S}p_{k}$ (13)
is the probability associated to the number of free predators, $n_{A}=n_{A}^{0}-\sum_{i=1}^{S}n_{i}$. To carry out the derivation, we use a detailed-balance condition satisfied by transition probabilities. The details can be found in A.
Second, using the generating function formalism, we find an analytical solution for out-of-equilibrium probabilities $P_{\bm{n}}(t)$ for the particular case of equal relaxation rates, $\nu_{1}=\dots=\nu_{S}$, and for an initial state formed only by searching predators. In the second case, we also show that the distribution is multinomial. From the form of the distributions (and the generating function), it is apparent that the model is perfectly identifiable out of equilibrium.
In the general case of different relaxation rates, we cannot analytically solve the master equation. However, it is a linear system of differential equations, so it could be easily integrated numerically. Besides numerical integration, there are other exact methods —based on matrix exponentiation (Keeling and Ross, 2009) or Laplace transforms (Etienne and Alonso, 2007)— or approximate —dimensionality reduction (Keeling and Ross, 2009)— that allow the computation of probabilities over time. Although the parameter estimation analysis presented below would be numerically feasible in the case of distinct relaxation rates, we do not discuss it in this contribution.
@@ 3.1 Out-of-equilibrium probability distribution
$\frac{\partial G}{\partial t}=\sum_{i=1}^{S}(1-z_{i})\biggl{(}\nu_{i}\frac{\partial G}{\partial z_{i}}+\theta_{i}\sum_{j=1}^{S}z_{j}\frac{\partial G}{\partial z_{j}}-\theta_{i}n_{A}^{0}G\biggr{)}.$
@@
We leave the details of the derivation of Eq. (14) from (10) to B.
@@
$K=\frac{\nu+R}{Q-R}=\frac{\nu+\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}}{\sum_{j=1}^{S}\theta_{j}(1-\hat{z}_{j})},$ (22)
where we have defined $R:=\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}$. Observe that the initial condition vector $\bm{\hat{z}}$ depends on the vector of $z$-variables, $\bm{z}$, via the explicit solution of the differential equations for $z_{i}$ in Eq. (17). Because the solution of the system of ODEs satisfied by the characteristic curves will depend on the vector $\bm{\hat{z}}$ of initial conditions, we have to substitute $\hat{z}_{j}$ as function of $\bm{z}$ at the end of the calculation in order to get the desired expression for the generating function $G(\bm{z},t)$.
@@
$Q-m(\bm{z})=\frac{KQ-\nu e^{(\nu+Q)t}+e^{\nu t}[(1+K)R+\nu-KQ]}{K+e^{(\nu+Q)t}},$
because $\sum_{j=1}^{S}\theta_{j}z_{j}=Q-m(\bf{z})$ by (15). Given that $\hat{z}_{j}$ are functions of $\bm{z}$, $K$ becomes itself a function of $\bm{z}$, which can be determined as follows. From Eq. (22) we can solve for $R=\sum_{j=1}^{S}\theta_{j}\hat{z}_{j}$ in terms of $K$ as $R=(KQ-\nu)/(1+K)$ which, upon substitution into the last expression, yields
$Q-m(\bm{z})=\frac{KQ-\nu e^{(\nu+Q)t}}{K+e^{(\nu+Q)t}}.$
In this equality, any dependence on the initial condition $\bm{\hat{z}}$ has been removed. We can solve for $K$ in terms of $m$, giving
$K=K(\bm{z},t)=\left(\frac{\nu+Q}{m(\bm{z})}-1\right)e^{(\nu+Q)t}.$ (25)
Finally, an explicit expression of initial conditions $\hat{z}_{j}$ as function of $\bm{z}$ is determined by substituting $K(\bm{z},t)$, given by Eq. (25), into (24), which results in an equation from which we can eliminate $\hat{z}_{j}$ in terms of $\bm{z}$. This manipulation gives
$\hat{z}_{j}=1-\left(1+\frac{\nu}{Q}\right)\frac{m(\bm{z})+[(1-z_{j})Q-m(\bm{z})]e^{Qt}}{m(\bm{z})+[\nu+Q-m(\bm{z})]e^{(\nu+Q)t}}.$ (26)
@@
$G(\bm{z},t)=\left[1+\frac{m({\bm{z}})(e^{-(\nu+Q)t}-1)}{\nu+Q}\right]^{n_{A}^{0}}\prod_{j=1}^{S}\left[1-\left(1+\frac{\nu}{Q}\right)\frac{m(\bm{z})+[Q(1-z_{j})-m(\bm{z})]e^{Qt}}{m(\bm{z})+[\nu+Q-m(\bm{z})]e^{(\nu+Q)t}}\right]^{\hat{n}_{j}}.$ (27)
Recall the definitions $m(\bm{z}):=\sum_{j=1}^{S}\theta_{j}(1-z_{j})$ and $Q:=\sum_{j=1}^{S}\theta_{j}$, and that the condition $n_{A}^{0}-\sum_{j=0}^{S}\hat{n}_{j}\geq 0$ must hold for the last expression to make sense. This function satisfies the initial condition (16) as well as the normalization condition $G(\bm{1},t)=1$. The generating function has to be expanded in power series of $z_{i}$ in order to obtain the explicit expression of the probability distribution $P_{\bm{n}}(t)$. Such task is hard in general. As shown by the derivation itself, it is very difficult to find the out-of-equilibrium distribution from the master equation, in contrast with the solution at stationarity: for the former, we cannot find it unless assuming equal handling rates, and we cannot carry out the series expansion to obtain the probability distribution $P_{\bm{n}}(t)$ for arbitrary initial conditions.
@@
$p_{0}(t)=\frac{1+e^{-\left(\nu+\sum_{k=1}^{S}\theta_{k}\right)t}\sum_{i=1}^{S}\frac{\theta_{i}}{\nu}}{1+\sum_{k=1}^{S}\frac{\theta_{k}}{\nu}},\qquad p_{j}(t)=\frac{\theta_{j}}{\nu}\frac{1-e^{-\left(\nu+\sum_{k=1}^{S}\theta_{k}\right)t}}{1+\sum_{k=1}^{S}\frac{\theta_{k}}{\nu}},$ (29)
for $j=1,\dots,S$. It is easy to check that $0\leq p_{j}(t)\leq 1$ for $j=0,1,\dots,S$ and arbitrary time $t$. In addition, $\sum_{j=0}^{S}p_{j}(t)=1$ holds for any time. Therefore, the parameter combinations $p_{j}(t)$ can be interpreted as probabilities. As we immediately show below, they are precisely the probabilities associated to the $j$-th population in a multinomial distribution.
@@
$G(\bm{z},t)=\biggl{(}p_{0}(t)+\sum_{j=1}^{S}p_{j}(t)z_{j}\biggr{)}^{n_{A}^{0}}=\sum_{n_{A}=0}^{n_{A}^{0}}\binom{n_{A}^{0}}{n_{A}}p_{0}^{n_{A}}\biggl{(}\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}-n_{A}}.$
@@
$\biggl{(}\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}-n_{A}}=\!\!\!\!\sum_{\begin{subarray}{c}n_{1},\dots,n_{S}\geq 0\\ \sum_{i}n_{i}=n_{A}^{0}-n_{A}\end{subarray}}\!\!\!\!\frac{(n_{A}^{0}-n_{A})!}{n_{1}!\cdots n_{S}!}\prod_{k=1}^{S}(p_{k}z_{k})^{n_{k}},$
@@
$P_{\bm{n}}(t)=\sum_{\begin{subarray}{c}n_{A}=0\\ n_{A}=n_{A}^{0}-\sum_{i}n_{i}\end{subarray}}^{n_{A}^{0}}\frac{n_{A}^{0}!}{n_{A}!n_{1}!\cdots n_{S}!}p_{0}^{n_{A}}p_{1}^{n_{1}}\cdots p_{S}^{n_{S}}$